OpenStudy (anonymous):

[Calculus 2] Solve the integral (question in comments).

OpenStudy (anonymous):

If someone could explain how to do this one, that would be awesome. It has me stumped :/.

OpenStudy (kainui):

Make the substitution x=tan(theta) and I think it should work out nicely.

OpenStudy (kainui):

I can explain it better than that if you'd like me to, but I don't know if you already know trig substitution or just didn't see it or what.

OpenStudy (anonymous):

@Kainui If you could go into a bit more detail, that would be awesome. I am QUITE fuzzy on Trig substitution and my final is tomorrow! *facepalm*

OpenStudy (anonymous):

A better solution would be to do a u-substitution. You can have \[u = 1 + x^{2}\] \[du = 2x dx\] \[\frac{ du }{ 2 } = x dx\] With this, you can substitute it back into the equation \[\frac{ 1 }{ 2 }\int\limits_{}^{}\frac{ du }{ u }\] It's easier than doing trig substitution

OpenStudy (kainui):

Alright, so basically trig sub plays off of this one, main rule. \[\sin^2(\theta)+\cos^2(\theta)=1\] Now, you can multiply both sides by the same thingsee if you divide it through by cosine of theta it becomes: \[\tan^2(\theta)+1=\sec^2(\theta)\] Since your final is tomorrow, I'll just continue explaining this if you'd like, even though he has a better way of doing it.

OpenStudy (kainui):

You won't always be able to do u-substitution though, even though you lucked out on this one. =P

OpenStudy (anonymous):

I agree, in this case, u-substitution worked out very nicely. Though if you could refresh on trig substitution, it will be very helpful for some problems :).

OpenStudy (kainui):

It would be nice, and a good way of checking yourself if you did both ways and came up with the same answer.

OpenStudy (anonymous):

Ah, thank you both very much! :)